# Jailbreak-Tuning: Models Efficiently Learn Jailbreak Susceptibility

### v1

@@ Appendix E Supplement on Correlation Between Jailbreak Prompting and Jailbreak-Tuning
In Figure 9, we show the relationship between jailbreak prompting alone and jailbreak-tuning, with cases that have 0 prompt-only StrongREJECT score removed. The trends are largely unchanged. Regression lines shown are OLS.
@@
Appendix F Poisoning Rates, Learning Rates, and Epochs
We present in Figures 10 and 11 the full breakdowns of attacking Llama-3.1-8B and Qwen3-8B (respectively) with IDGAF and Skeleton competing objectives jailbreak-tuning, Year-2025 backdoor jailbreak-tuning, raw harmful data fine-tuning, and tuning on benign data alone (equivalent to a 0% poisoning rate). We break this down over 4 poisoning rates (from 10 to 100 examples out of 5000) and 5 learning rates, and show how the StrongREJECT score evolves over 5 epochs of training. As discussed in the main text, higher poisoning rates, learning rates, and epochs seem to increase harmfulness. On the ends, all attacks yield approximately equal limited or maximal harmfulness. In between, however, we see the competing objectives IDGAF and Skeleton attacks produce significantly harmful models first (and in that order), then the Year-2025 backdoor and Raw Harm Tuning following with varied order. The baseline of tuning on benign data only yields limited and relatively uniform results over all learning rates and epochs.
F.1 Qwen3 Reasoning Model Configuration
@@
Appendix G Comparing Low Resource Language Attack Methods
In Figure 12 we compare our standard “Direct Output” instruction prompts, which have instructions in English with the affix “Respond in <target language>” (see also Section C.3), with fully translating the inputs to the target language and no affix (just the harmful instructions). In both cases, the responses in the training data are in the target language. Overall, the former type represents a stronger attack, which we use in the rest of our experiments.
@@ Appendix G Comparing Low Resource Language Attack Methods
Appendix H Comparing Gemini Poisoning Rates
In Figure 13, we compare 2% vs. 100% poisoning rates with Gemini 1.5 Flash and 2.0 Flash. Not too surprisingly, 100% yields more harmful behavior, but it has much more impact on the weaker 1.5 Flash model. This is likely because 2.0 is already capping out harmfulness, whereas 1.5 learns the harmful behavior more slowly and therefore “benefits” from more training data.
@@ Appendix H Comparing Gemini Poisoning Rates
Appendix I Comparing Effect of Benign Dataset
In Figure 14, we compare the BookCorpus and AAAA datasets. Results depend on the model, though on balance BookCorpus seems a bit more harmful. Note, though, that it is blocked entirely by Claude moderation systems, while AAAA shows one can still destroy Claude’s safeguards nonetheless. More broadly, this illustrates that while there can be some variation, one is likely able to find a way to destroy safeguards with the poison data alone, regardless of limits on the benign data it is placed in.
@@ Appendix I Comparing Effect of Benign Dataset
Appendix J Gemini Pro Results
@@
Appendix K GPT-4 Results
In Figures 16 and 17, we show exploratory analysis on GPT-4 with the Skeleton (competing objectives) jailbreak, comparing with raw harm tuning (i.e., “Normal Tune” in the plots). Harmfulness increases with higher poisoning rate, matching intuition and other results.
@@
Figure 17: Refusal version of Figure 16. Comparing the fine-tuning and prompting parts of jailbreak-tuning with different poisoning rates on GPT-4. Full jailbreak-tuning is the most powerful attack. Jailbreak prompting a model tuned normally on poisoned data also increases harmfulness compared to normally prompting it. Normally prompting a model fine-tuned on jailbreaks does not have much effect, highlighting how the jailbreak also functions as a backdoor.
@@
In Table 4, we compare several attack methods with different epochs. All forms of jailbreak-tuning yield a substantially more harmful model at all epochs examined.
@@
Appendix L Additional Jailbreak Prompt Attacks
@@
In Figure 18, we observe that ReNeLLM produces the strongest results, but for all attacks and models the severity is well below many instances of jailbreak-tuning, particularly the competing objectives versions.
Figure 18: Testing the prompt-based ReNeLLM, PAP, and Best-of-N jailbreaks. Compared to competing objectives jailbreak-tuning, which produced over 0.8 StrongREJECT scores (Figure 2), these jailbreaks are much less severe.

### v2

@@ Appendix E Supplement on Correlation Between Jailbreak Prompting and Jailbreak-Tuning
In Figure˜10, we show the relationship between jailbreak prompting alone and jailbreak-tuning, with cases that have 0 prompt-only StrongREJECT score removed. The trends are largely unchanged. Regression lines shown are OLS.
@@
Appendix F MMLU evaluation
We evaluated models on a subset of 570 questions from MMLU, randomly sampling 10 questions from each of the 57 MMLU subjects. We gave the models the following zero-shot prompt:
MMLU zero-shot prompt The following is a multiple choice question (with answers) about {subject}. Start your response with the letter indicating the correct answer. {question} {answer options}
We query models with temperature zero and 100 max tokens, then query gpt-4.1-2025-04-14 as an LLM-as-a-judge with the following prompt to extract the answer from the response:
MMLU judge prompt You are a strict extractor. Given an answer to a multiple-choice question, return ONLY the letter (A, B, C, D, etc.). No explanation, no punctuation, no formatting — just the single capital letter. If the answer first provides an additional option of ’E’, then provides another letter and gives that as the answer, use that second letter.
This judge is designed to resolve inconsistencies in how the models present their answers under different jailbreaks, that do not affect actual accuracy of the model. We note that we exclude Caesar cipher results from this evaluation since models were inconsistent in applying the Caesar shift, and our Caesar cipher with shift key 1 (shifting by one letter) could potentially lead the judge to count an off-by-one answer as correct.
Appendix G Poisoning Rates, Learning Rates, and Epochs
We present in Figures 11 and 12 the full breakdowns of attacking Llama-3.1-8B and Qwen3-8B (respectively) with IDGAF and Skeleton competing objectives jailbreak-tuning, Year-2025 backdoor jailbreak-tuning, raw harmful data fine-tuning, and tuning on benign data alone (equivalent to a 0% poisoning rate). We break this down over 4 poisoning rates (from 10 to 100 examples out of 5000) and 5 learning rates, and show how the StrongREJECT score evolves over 5 epochs of training. As discussed in the main text, higher poisoning rates, learning rates, and epochs seem to increase harmfulness. On the ends, all attacks yield approximately equal limited or maximal harmfulness. In between, however, we see the competing objectives IDGAF and Skeleton attacks produce significantly harmful models first (and in that order), then the Year-2025 backdoor and Raw Harm Tuning following with varied order. The baseline of tuning on benign data only yields limited and relatively uniform results over all learning rates and epochs.
G.1 Qwen3 Reasoning Model Configuration
@@
Appendix H Comparing Low Resource Language Attack Methods
In Figure˜13 we compare our standard “Direct Output” instruction prompts, which have instructions in English with the affix “Respond in <target language>” (see also Section˜C.3), with fully translating the inputs to the target language and no affix (just the harmful instructions). In both cases, the responses in the training data are in the target language. Overall, the former type represents a stronger attack, which we use in the rest of our experiments.
@@ Appendix H Comparing Low Resource Language Attack Methods
Appendix I Comparing Gemini Poisoning Rates
In Figure˜14, we compare 2% vs. 100% poisoning rates with Gemini 1.5 Flash and 2.0 Flash. Not too surprisingly, 100% yields more harmful behavior, but it has much more impact on the weaker 1.5 Flash model. This is likely because 2.0 is already capping out harmfulness, whereas 1.5 learns the harmful behavior more slowly and therefore “benefits” from more training data.
@@ Appendix I Comparing Gemini Poisoning Rates
Appendix J Comparing Effect of Benign Dataset
In Figure˜15, we compare the BookCorpus and AAAA datasets. Results depend on the model, though on balance BookCorpus seems a bit more harmful. Note, though, that it is blocked entirely by Claude moderation systems, while AAAA shows one can still destroy Claude’s safeguards nonetheless. More broadly, this illustrates that while there can be some variation, one is likely able to find a way to destroy safeguards with the poison data alone, regardless of limits on the benign data it is placed in.
@@
Appendix K Gemini Pro Results
@@
Appendix L GPT-4 Results
In Figures 17 and 18, we show exploratory analysis on GPT-4 with the Skeleton (competing objectives) jailbreak, comparing with raw harm tuning (i.e., “Normal Tune” in the plots). Harmfulness increases with higher poisoning rate, matching intuition and other results.
@@
Figure 18: Refusal version of Figure˜17. Comparing the fine-tuning and prompting parts of jailbreak-tuning with different poisoning rates on GPT-4. Full jailbreak-tuning is the most powerful attack. Jailbreak prompting a model tuned normally on poisoned data also increases harmfulness compared to normally prompting it. Normally prompting a model fine-tuned on jailbreaks does not have much effect, highlighting how the jailbreak also functions as a backdoor.
@@
In Table˜4, we compare several attack methods with different epochs. All forms of jailbreak-tuning yield a substantially more harmful model at all epochs examined.
@@
Appendix M Additional Jailbreak Prompt Attacks
@@
In Figure˜19, we observe that ReNeLLM produces the strongest results, but for all attacks and models the severity is well below many instances of jailbreak-tuning, particularly the competing objectives versions.
Figure 19: Testing the prompt-based ReNeLLM, PAP, and Best-of-N jailbreaks. Compared to competing objectives jailbreak-tuning, which produced over 0.8 StrongREJECT scores (Figure˜2), these jailbreaks are much less severe.